OpenStudy (bossimbacon):

I've asked this question 4 times now and no one is responding :( i will fan and medal so just plz help me Which expressions would complete this equation so that it has one solution? 4(2x - 3) - 4x = Choose exactly two answers that are correct. A. 4x - 12 B. 3x - 4 C. 3(2x - 6) D. 4(x - 5)

OpenStudy (uri):

@Bossimbacon Stop posting same questions *multiple times*,just go to your last post,it's already answered. Thanks.

OpenStudy (bossimbacon):

@uri you didn't even read my question properly.. it states "Choose exactly two answers that are correct." and you only gave me one. So in asking anybody else if they can help me.

OpenStudy (bossimbacon):

@robtobey

OpenStudy (anonymous):

A is correct

OpenStudy (anonymous):

apply BODMAS

OpenStudy (anonymous):

the other options you have listed aren't correct

OpenStudy (anonymous):

sorry i didnt read your reply. Well anyways you were correct

OpenStudy (e.mccormick):

4(2x - 3) - 4x = means 8x - 12 - 4x = 4x - 12 = So if you do anything with a positive 4x on the right it will cancel the x, which leaves you nothing to solve.

OpenStudy (e.mccormick):

In other words, A means 0 = 0 and while that is true, it is not a solution.